OpenStudy (anonymous):

What is the value of y given that logx 1 = y(x>0) ?

OpenStudy (kropot72):

\[\log_{x} 1=y\] \[x ^{y}=1\] \[y \times \log_{10} x=\log_{10} 1\] \[y=\frac {\log_{10} 1}{\log_{10} x}\] Can you finish it now?

OpenStudy (experimentx):

\[ x ^{y}=1 = x^0 \\ \\ \\ and, \; \log_{k} 1 = 0 \] k may have any value!!

OpenStudy (kropot72):

By definition the log of 1 to any base must equal zero.

OpenStudy (anonymous):

Any number to the zeroth power gives you 1. Here is the best reasoing I found: